# Dual Quaternion SE(3) Synchronization with Recovery Guarantees

### v1

@@ Appendix C Proofs in Section 2
Case 2: $a_{k}=0$. Then $s_{k}=b_{k}\epsilon$ has zero standard part, so the objective $z^{*}s_{k}+(z^{*}s_{k})^{*}$ has zero standard part for every $z\in\mathrm{UDQ}$. We may thus adopt an arbitrary tie-breaking rule to select an optimizer $z_{k}$. In this case $p_{k}=z_{k}^{*}s_{k}$ is purely infinitesimal (i.e., $\operatorname{st}(p_{k})=0$), and the identity above still yields
@@
D.5 Proof of Corollary 3.4
@@
As $k\to\infty$, using $|\lambda_{1,_{st}}|>|\lambda_{2,_{st}}|$ and standard dual-number expansion,
@@
$\left(\sqrt{\sum_{j=1}^{n}|\lambda_{j}^{k}\alpha_{j}|^{2}}-|\lambda_{1}|^{k}|\alpha_{1}|\right)\left(\sqrt{\sum_{j=1}^{n}|\lambda_{j}^{k}\alpha_{j}|^{2}}+|\lambda_{1}|^{k}|\alpha_{1}|\right)\left(\sqrt{\sum_{j=1}^{n}|\lambda_{j}^{k}\alpha_{j}|^{2}}+|\lambda_{1}|^{k}|\alpha_{1}|\right)^{-1}$
@@
$=\sum_{j=2}^{n}|\lambda_{j,st}|^{2k}(1+2k\lambda_{j,st}^{-1}\lambda_{j,\mathcal{I}}\epsilon)|\alpha_{j}|^{2}A^{-1}$
$=O_{D}\left(\left(\frac{{\lambda_{2,st}}^{2}}{\lambda_{1,st}}\right)^{k}\right),$
@@
$A:=\Big(\sum_{j=1}^{n}|\lambda_{j,st}|^{2k}(1+2k\lambda_{j,st}^{-1}\lambda_{j,\mathcal{I}}\epsilon)|\alpha_{j}|^{2}\Big)^{1/2}+|\lambda_{1,st}|^{k}(1+k\lambda_{1,st}^{-1}\lambda_{1,\mathcal{I}}\epsilon)|\alpha_{1}|.$
@@
$t_{j}^{k}=\lambda_{j}^{k}\alpha_{j}\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big]^{-1}.$
@@
$t_{1}^{k}-\lambda_{1}^{k}\alpha_{1}\left(|\lambda_{1}|^{k}|\alpha_{1}|\right)^{-1}$ $=\lambda_{1}^{k}\alpha_{1}\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big]^{-1}-\lambda_{1}^{k}\alpha_{1}\left(|\lambda_{1}|^{k}|\alpha_{1}|\right)^{-1}$
$=-\lambda_{1}^{k}\alpha_{1}\left(|\lambda_{1}|^{k}|\alpha_{1}|\right)^{-1}\cdot O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big[\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big]^{-1}$
$=-\alpha_{1}|\alpha_{1}|^{-1}\cdot O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big]^{-1}.$
@@
$\bm{w}^{k}$ $=\bm{v}_{1}\alpha_{1}|\alpha_{1}|^{-1}\Bigg\{1-O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big]^{-1}\Bigg\}$
$+\sum_{j=2}^{n}\bm{v}_{j}\lambda_{j}^{k}\alpha_{j}\Big(|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big)^{-1}.$

### v2

@@ Appendix C Proofs in Section 2
We first prove the estimate componentwise. Fix $i\in\{1,\ldots,n\}$, and write $r_{i}=q_{i}+\epsilon p_{i}$, $h_{i}=e_{i}+\epsilon f_{i}$, where $q_{i},p_{i},e_{i},f_{i}\in\mathbb{H}$. Since $r_{i}\in{\rm UDQ}$, we have
$|q_{i}|=1,\qquad{\rm sc}(q_{i}^{*}p_{i})=0.$
Let $\Pi(r_{i}+h_{i})=\tilde{q}_{i}+\epsilon\tilde{p}_{i}$. The local-chart condition (22) gives
$|q_{i}+e_{i}|\geq\frac{1}{2},$
in particular $q_{i}+e_{i}\neq 0$. By the definition in Proposition 2.7, we therefore have
$\tilde{q}_{i}=\frac{q_{i}+e_{i}}{|q_{i}+e_{i}|}.$
For any unit quaternion $q$, define the tangent projection
$P_{q}(\xi):=\xi-q\,{\rm sc}(q^{*}\xi),\qquad\xi\in\mathbb{H}.$
Then $P_{q}$ is the Euclidean orthogonal projection onto
$T_{q}\mathbb{S}^{3}=\{\xi\in\mathbb{H}:{\rm sc}(q^{*}\xi)=0\}.$
In particular, $\|P_{q}\|\leq 1$ and $P_{q_{i}}(p_{i})=p_{i}$. Using the explicit form of the dual part of the componentwise normalization, we have
$\tilde{p}_{i}=\frac{1}{|q_{i}+e_{i}|}P_{\tilde{q}_{i}}(p_{i}+f_{i}).$
Subtracting $p_{i}=P_{q_{i}}(p_{i})$, we decompose the difference as
$\tilde{p}_{i}-p_{i}$ $=\frac{1}{|q_{i}+e_{i}|}P_{\tilde{q}_{i}}(f_{i})+\left(\frac{1}{|q_{i}+e_{i}|}-1\right)P_{\tilde{q}_{i}}(p_{i})+\bigl(P_{\tilde{q}_{i}}-P_{q_{i}}\bigr)p_{i}.$
We bound these three terms separately.
First, since $|q_{i}+e_{i}|\geq 1/2$ and $\|P_{\tilde{q}_{i}}\|\leq 1$,
$\left|\frac{1}{|q_{i}+e_{i}|}P_{\tilde{q}_{i}}(f_{i})\right|\leq 2|f_{i}|.$
Second, again using $|q_{i}+e_{i}|\geq 1/2$, together with the reverse triangle inequality, we obtain
$\left|\frac{1}{|q_{i}+e_{i}|}-1\right|=\frac{\bigl||q_{i}+e_{i}|-1\bigr|}{|q_{i}+e_{i}|}\leq 2|e_{i}|,$
and therefore
$\left|\left(\frac{1}{|q_{i}+e_{i}|}-1\right)P_{\tilde{q}_{i}}(p_{i})\right|\leq 2|p_{i}|\,|e_{i}|.$
It remains to control the change of tangent projection from $P_{q_{i}}$ to $P_{\tilde{q}_{i}}$. For any $\xi\in\mathbb{H}$, we write
$\bigl(P_{\tilde{q}_{i}}-P_{q_{i}}\bigr)\xi$ $=q_{i}\,{\rm sc}(q_{i}^{*}\xi)-\tilde{q}_{i}\,{\rm sc}(\tilde{q}_{i}^{*}\xi)=(q_{i}-\tilde{q}_{i})\,{\rm sc}(q_{i}^{*}\xi)+\tilde{q}_{i}\,{\rm sc}\bigl((q_{i}-\tilde{q}_{i})^{*}\xi\bigr).$
Using
$|{\rm sc}(q_{i}^{*}\xi)|\leq|\xi|,\qquad|{\rm sc}((q_{i}-\tilde{q}_{i})^{*}\xi)|\leq|q_{i}-\tilde{q}_{i}|\,|\xi|,\qquad|\tilde{q}_{i}|=1,$
we obtain $|(P_{\tilde{q}_{i}}-P_{q_{i}})\xi|\leq 2|\tilde{q}_{i}-q_{i}|\,|\xi|,$ and then
$\|P_{\tilde{q}_{i}}-P_{q_{i}}\|\leq 2|\tilde{q}_{i}-q_{i}|\leq 4|e_{i}|.$
Consequently,
$\left|\bigl(P_{\tilde{q}_{i}}-P_{q_{i}}\bigr)p_{i}\right|\leq 4|p_{i}|\,|e_{i}|.$
Combining the three estimates above gives the following elementwise bound for the dual part:
$\left|\bigl(\Pi(r_{i}+h_{i})-r_{i}\bigr)_{\mathcal{I}}\right|=|\tilde{p}_{i}-p_{i}|\leq 2|f_{i}|+6|p_{i}|\,|e_{i}|.$
We now pass from the elementwise estimates to the vectorized estimates. Since $\Pi$ acts componentwise on $\bm{x}\in\mathbb{DH}^{n}$ when $x_{i,\mathrm{st}}\neq 0$, we have
$\Pi(\bm{r}+\bm{h})_{i}=\Pi(r_{i}+h_{i}).$
For the dual part, we use the elementwise dual estimate and the triangle inequality in $\ell_{2}$. This yields
$\left\|\bigl(\Pi(\bm{r}+\bm{h})-\bm{r}\bigr)_{\mathcal{I}}\right\|_{2}$ $\leq 2\left(\sum_{i=1}^{n}|(h_{i})_{\mathcal{I}}|^{2}\right)^{1/2}+6\left(\sum_{i=1}^{n}|(r_{i})_{\mathcal{I}}|^{2}|(h_{i})_{\rm st}|^{2}\right)^{1/2}$
$\leq 2\|\bm{h}_{\mathcal{I}}\|_{2}+6\max_{1\leq i\leq n}|(r_{i})_{\mathcal{I}}|\,\|\bm{h}_{\rm st}\|_{2}.$
This proves (23).
D.5 Proof of Lemma D.5
@@
Case 2: $a_{k}=0$. Then $s_{k}=b_{k}\epsilon$. If $b_{k}=0$, then $p_{k}=0$ for every $z_{k}\in{\rm UDQ}$, and
@@
Assume now that $b_{k}\neq 0$. Write $z_{k}^{*}=c_{k}+d_{k}\epsilon$. Then
$p_{k}=z_{k}^{*}s_{k}=c_{k}b_{k}\epsilon,\qquad p_{k}+p_{k}^{*}=2\,{\rm sc}(c_{k}b_{k})\epsilon.$
Since the standard part of the objective is zero for every unit $c_{k}$, we break ties lexicographically by maximizing ${\rm sc}(c_{k}b_{k})$. This is achieved by choosing $c_{k}=\frac{b_{k}^{*}}{|b_{k}|}$ and $d_{k}=0$. Hence $p_{k}=|b_{k}|\epsilon\in\mathbb{D}$, so $p_{k}=p_{k}^{*}$ and $0\preceq p_{k}\preceq n+0\epsilon$. Moreover,
$\frac{1}{2}\operatorname{d}(\bm{x}^{k},\hat{\bm{x}})^{2}=n-\frac{p_{k}+p_{k}^{*}}{2}=n-|b_{k}|\epsilon=|p_{k}-n|.$
@@
D.6 Proof of Lemma D.6
By the scale invariance of the componentwise normalization map $\Pi$,
$\bm{x}^{k+1}=\Pi(\bm{C}\bm{x}^{k})=\Pi\left(\frac{1}{n}\bm{C}\bm{x}^{k}\right).$
Since $\bm{C}-\bm{\Delta}=\hat{\bm{x}}\hat{\bm{x}}^{*}$, we write
$\frac{1}{n}\bm{C}\bm{x}^{k}=\bm{r}^{k}+\bm{h}^{k},$
where
$\bm{h}^{k}:=\frac{1}{n}\bigl((\bm{C}-\bm{\Delta})\bm{x}^{k}-n\bm{r}^{k}\bigr)+\frac{1}{n}\bm{\Delta}(\bm{x}^{k}-\bm{r}^{k})+\frac{1}{n}\bm{\Delta}\bm{r}^{k}.$
By Assumption (8), the normalization input satisfies
$\min_{i}\left|\left(\bm{r}_{i}^{k}+\bm{h}_{i}^{k}\right)_{\rm st}\right|=\min_{i}\left|\left(\frac{1}{n}\bm{C}\bm{x}^{k}\right)_{i,{\rm st}}\right|\geq\frac{1}{2}.$
Therefore, applying Lemma D.4 with $\bm{r}=\bm{r}^{k}$ and $\bm{h}=\bm{h}^{k}$, and using Assumption (7), we get
$\left\|\bigl(\bm{x}^{k+1}-\bm{r}^{k}\bigr)_{\mathcal{I}}\right\|_{2}\leq 2\|(\bm{h}^{k})_{\mathcal{I}}\|_{2}+6B\|(\bm{h}^{k})_{\rm st}\|_{2},$
and
$\left\|\bigl(\bm{x}^{k+1}-\bm{r}^{k}\bigr)_{\rm st}\right\|_{2}\leq 2\|(\bm{h}^{k})_{\rm st}\|_{2}.$
Then the stable gauge-switching condition (9) gives
$\delta_{\mathcal{I}}^{k+1}=\left\|\bigl(\bm{x}^{k+1}-\hat{\bm{x}}z_{k+1}\bigr)_{\mathcal{I}}\right\|_{2}$ $\leq\left\|\bigl(\bm{x}^{k+1}-\hat{\bm{x}}z_{k}\bigr)_{\mathcal{I}}\right\|_{2}+\gamma\left\|\bigl(\bm{x}^{k+1}-\hat{\bm{x}}z_{k}\bigr)_{\rm st}\right\|_{2}$
$\leq 2\|(\bm{h}^{k})_{\mathcal{I}}\|_{2}+(6B+2\gamma)\|(\bm{h}^{k})_{\rm st}\|_{2}.$
It remains to bound the standard and dual components of $\bm{h}^{k}$. Define the scalar dual number $\eta^{k}:=z_{k}^{*}\hat{\bm{x}}^{*}\bm{x}^{k}-n$. Then
$(\bm{C}-\bm{\Delta})\bm{x}^{k}-n\bm{r}^{k}=\hat{\bm{x}}\hat{\bm{x}}^{*}\bm{x}^{k}-n\hat{\bm{x}}z_{k}=\bm{r}^{k}\eta^{k}.$
By Lemma D.5, we have
$|\eta^{k}|=\frac{1}{2}\operatorname{d}(\bm{x}^{k},\hat{\bm{x}})^{2}.$
Since this is an equality of dual numbers, its standard part gives
$|(\eta^{k})_{\rm st}|=\frac{1}{2}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})^{2}.$
Its infinitesimal component satisfies
[…]
Moreover, for any $\bm{u}=\bm{u}_{\rm st}+\epsilon\bm{u}_{\mathcal{I}}$, the infinitesimal coefficient of the dual-valued norm satisfies $|(\|\bm{u}\|_{2})_{\mathcal{I}}|\leq\|\bm{u}_{\mathcal{I}}\|_{2}$. Applying this to $\bm{u}=\bm{x}^{k}-\bm{r}^{k}$ gives $|\operatorname{d}_{\mathcal{I}}(\bm{x}^{k},\hat{\bm{x}})|\leq\delta_{\mathcal{I}}^{k}$. Hence
$|(\eta^{k})_{\mathcal{I}}|\leq\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})\delta_{\mathcal{I}}^{k}.$ (33)
Now, we first bound $\|(\bm{h}^{k})_{\rm st}\|_{2}$. From the definition of $\bm{h}^{k}$,
$\|(\bm{h}^{k})_{\rm st}\|_{2}$ $\leq\frac{1}{n}\left\|\bigl((\bm{C}-\bm{\Delta})\bm{x}^{k}-n\bm{r}^{k}\bigr)_{\rm st}\right\|_{2}+\frac{1}{n}\left\|\bigl(\bm{\Delta}(\bm{x}^{k}-\bm{r}^{k})\bigr)_{\rm st}\right\|_{2}+\frac{1}{n}\left\|(\bm{\Delta}\bm{r}^{k})_{\rm st}\right\|_{2}.$
For the first term, since $(\bm{C}-\bm{\Delta})\bm{x}^{k}-n\bm{r}^{k}=\bm{r}^{k}\eta^{k}$, we have
$\left\|\bigl((\bm{C}-\bm{\Delta})\bm{x}^{k}-n\bm{r}^{k}\bigr)_{\rm st}\right\|_{2}=\|(\bm{r}^{k})_{\rm st}\|_{2}|(\eta^{k})_{\rm st}|=\frac{\sqrt{n}}{2}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})^{2}.$
For the second term,
[…]
For the third term, using $\bm{r}^{k}=\hat{\bm{x}}z_{k}$ and the fact that right multiplication by the unit quaternion $(z_{k})_{\rm st}$ preserves the Euclidean norm,
$\left\|(\bm{\Delta}\bm{r}^{k})_{\rm st}\right\|_{2}=\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}.$
Therefore,
$\|(\bm{h}^{k})_{\rm st}\|_{2}\leq\frac{1}{2\sqrt{n}}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})^{2}+\frac{1}{n}\|\bm{\Delta}\|_{{\rm op},{\rm st}}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})+\frac{1}{n}\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}.$
Next we bound $\|(\bm{h}^{k})_{\mathcal{I}}\|_{2}$. Again using $(\bm{C}-\bm{\Delta})\bm{x}^{k}-n\bm{r}^{k}=\bm{r}^{k}\eta^{k}$, we obtain
$\left\|\bigl((\bm{C}-\bm{\Delta})\bm{x}^{k}-n\bm{r}^{k}\bigr)_{\mathcal{I}}\right\|_{2}$ $=\|(\bm{r}^{k}\eta^{k})_{\mathcal{I}}\|_{2}\leq\|(\bm{r}^{k})_{\rm st}\|_{2}|(\eta^{k})_{\mathcal{I}}|+\|(\bm{r}^{k})_{\mathcal{I}}\|_{2}|(\eta^{k})_{\rm st}|.$
Since
$\|(\bm{r}^{k})_{\rm st}\|_{2}=\sqrt{n},\qquad\|(\bm{r}^{k})_{\mathcal{I}}\|_{2}\leq B\sqrt{n},$ (34)
this together with (33) gives
$\left\|\bigl((\bm{C}-\bm{\Delta})\bm{x}^{k}-n\bm{r}^{k}\bigr)_{\mathcal{I}}\right\|_{2}\leq\sqrt{n}\,\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})\delta_{\mathcal{I}}^{k}+\frac{B\sqrt{n}}{2}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})^{2}.$
For the perturbation term,
$\left\|\bigl(\bm{\Delta}(\bm{x}^{k}-\bm{r}^{k})\bigr)_{\mathcal{I}}\right\|_{2}$ $\leq\|\bm{\Delta}\|_{{\rm op},{\rm st}}\left\|(\bm{x}^{k}-\bm{r}^{k})_{\mathcal{I}}\right\|_{2}+\|\bm{\Delta}\|_{{\rm op},\mathcal{I}}\left\|(\bm{x}^{k}-\bm{r}^{k})_{\rm st}\right\|_{2}$
[…]
Finally, from (34) we get
$\left\|(\bm{\Delta}\bm{r}^{k})_{\mathcal{I}}\right\|_{2}\leq\sqrt{n}\left(\|\bm{\Delta}\|_{{\rm op},\mathcal{I}}+B\|\bm{\Delta}\|_{{\rm op},{\rm st}}\right).$
Combining the three estimates yields
$\|(\bm{h}^{k})_{\mathcal{I}}\|_{2}$ $\leq\frac{1}{\sqrt{n}}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})\delta_{\mathcal{I}}^{k}+\frac{B}{2\sqrt{n}}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})^{2}+\frac{1}{n}\|\bm{\Delta}\|_{{\rm op},{\rm st}}\delta_{\mathcal{I}}^{k}$
$\quad+\frac{1}{n}\|\bm{\Delta}\|_{{\rm op},\mathcal{I}}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})+\frac{1}{\sqrt{n}}\left(\|\bm{\Delta}\|_{{\rm op},\mathcal{I}}+B\|\bm{\Delta}\|_{{\rm op},{\rm st}}\right).$
Substituting the bounds for $\|(\bm{h}^{k})_{\rm st}\|_{2}$ and $\|(\bm{h}^{k})_{\mathcal{I}}\|_{2}$ into $\delta_{\mathcal{I}}^{k+1}\leq 2\|(\bm{h}^{k})_{\mathcal{I}}\|_{2}+(6B+2\gamma)\|(\bm{h}^{k})_{\rm st}\|_{2}$ gives
$\delta_{\mathcal{I}}^{k+1}$ $\leq\left(\frac{2}{\sqrt{n}}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})+\frac{2}{n}\|\bm{\Delta}\|_{{\rm op},{\rm st}}\right)\delta_{\mathcal{I}}^{k}$
[…]
$\quad+\frac{2}{\sqrt{n}}\left(\|\bm{\Delta}\|_{{\rm op},\mathcal{I}}+B\|\bm{\Delta}\|_{{\rm op},{\rm st}}\right)+\frac{6B+2\gamma}{n}\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}.$
This is exactly (25).
D.7 Proof of Corollary 3.4
@@
As $k\to\infty$, using $|\lambda_{1,_{\mathrm{st}}}|>|\lambda_{2,_{\mathrm{st}}}|$ and standard dual-number expansion,
@@
$=\left(\sqrt{\sum_{j=1}^{n}|\lambda_{j}^{k}\alpha_{j}|^{2}}-|\lambda_{1}|^{k}|\alpha_{1}|\right)\left(\sqrt{\sum_{j=1}^{n}|\lambda_{j}^{k}\alpha_{j}|^{2}}+|\lambda_{1}|^{k}|\alpha_{1}|\right)\left(\sqrt{\sum_{j=1}^{n}|\lambda_{j}^{k}\alpha_{j}|^{2}}+|\lambda_{1}|^{k}|\alpha_{1}|\right)^{-1}$
@@
$=\sum_{j=2}^{n}|\lambda_{j,\mathrm{st}}|^{2k}(1+2k\lambda_{j,\mathrm{st}}^{-1}\lambda_{j,\mathcal{I}}\epsilon)|\alpha_{j}|^{2}A^{-1}$
$=O_{D}\left(\left(\frac{{\lambda_{2,\mathrm{st}}}^{2}}{\lambda_{1,\mathrm{st}}}\right)^{k}\right),$
@@
$A:=\Big(\sum_{j=1}^{n}|\lambda_{j,\mathrm{st}}|^{2k}(1+2k\lambda_{j,\mathrm{st}}^{-1}\lambda_{j,\mathcal{I}}\epsilon)|\alpha_{j}|^{2}\Big)^{1/2}+|\lambda_{1,\mathrm{st}}|^{k}(1+k\lambda_{1,\mathrm{st}}^{-1}\lambda_{1,\mathcal{I}}\epsilon)|\alpha_{1}|.$
@@
$t_{j}^{k}=\lambda_{j}^{k}\alpha_{j}\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big]^{-1}.$
@@
$t_{1}^{k}-\lambda_{1}^{k}\alpha_{1}\left(|\lambda_{1}|^{k}|\alpha_{1}|\right)^{-1}$ $=\lambda_{1}^{k}\alpha_{1}\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big]^{-1}-\lambda_{1}^{k}\alpha_{1}\left(|\lambda_{1}|^{k}|\alpha_{1}|\right)^{-1}$
$=-\lambda_{1}^{k}\alpha_{1}\left(|\lambda_{1}|^{k}|\alpha_{1}|\right)^{-1}\cdot O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big[\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big]^{-1}$
$=-\alpha_{1}|\alpha_{1}|^{-1}\cdot O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big]^{-1}.$
@@
$\bm{w}^{k}$ $=\bm{v}_{1}\alpha_{1}|\alpha_{1}|^{-1}\Bigg\{1-O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big]^{-1}\Bigg\}$
$\quad+\sum_{j=2}^{n}\bm{v}_{j}\lambda_{j}^{k}\alpha_{j}\Big(|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big)^{-1}.$